Expected Cash Flows

Year Project A Project B

0 -400 -575

1 95 150

2 110 200

3 118 250

4 125 275

5 140 230

6 150 180

a. If you were told that each project’s cost of capital was 10%, which project should be selected using the NPV criteria? b. What is each project’s IRR? c. What is the regular payback period for these two projects? d. What is the profitability index for each project if the cost of capital is 12%?


1
Expert's answer
2021-05-25T11:16:16-0400

a. "NPVa = -400+95\/1.1+110\/1.1^2+118\/1.1^3+125\/1.1^4+140\/1.1^5+150\/1.1^6 = 122.9."

"NPVb = -575+150\/1.1+200\/1.1^2+250\/1.1^3+275\/1.1^4+230\/1.1^5+180\/1.1^6 = 346.73."

b. IRR is a cost of capital, at which NPV = 0. IRR is higher for project B.

c. Payback period is shorter for B.

d. Profitability index is higher for B.